OpenStudy (anonymous):

A small rock sits on the edge of a tall building. A strong wind blows the rock off the edge. The distance, in feet, between the rock and the ground tseconds after the rock leaves the edge is given by d=−16t^2−4t+472. How many seconds after the rock leaves the edge is it 460 feet from the ground? How many seconds after the rock leaves the edge does it hit the ground? Please help!

OpenStudy (anonymous):

i'll fan and give a medal!

geerky42 (geerky42):

So \(d\) is distance between rock and ground. "How many seconds after the rock leaves the edge is it 460 feet from the ground?" So \(d=460\), so subtract both sides by 460 then apply quadratic formula. Just remember that t has to be positive. "How many seconds after the rock leaves the edge does it hit the ground? " This time, \(d=0\) So just apply quadratic formula right there. Again, t has to be positive.

OpenStudy (anonymous):

do you know the answers? i don't know what i'm doing wrong..

OpenStudy (anonymous):

@geerky42 for "How many seconds after the rock leaves the edge is it 460 feet from the ground?" I am getting 313.94

OpenStudy (anonymous):

Nevermind, I guess i put it in my calculator wrong.. I ended up getting .75 for the first answer. Now I am working on the second answer.

geerky42 (geerky42):

You know quadratic formula? \(t = \dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\), from \(0 = at^2+bt+c\)

geerky42 (geerky42):

Yeah I got that as well.

OpenStudy (anonymous):

How should I be putting this into my calculator? Everytime I try I get the wrong answer.

OpenStudy (anonymous):

@geerky42